OpenStudy (anonymous):

Thanks for helping! This question is really weird :( r=-6sin theta I need to multiply both sides of the equation by r and use r^2=x^2+y^2 to rewrite the equation in terms of x and y.

OpenStudy (anonymous):

@Ashleyisakitty @jim_thompson5910 @zepdrix @Nnesha @e.mccormick @wio @sammixboo @kropot72 @mathmate

OpenStudy (astrophysics):

Ok so you want to convert it to rectangular, it's good you notice we have to multiply r and use \[r^2 = x^2+y^2\] we know the ratio for sin theta is the following \[\sin \theta = \frac{ y }{ r }\] so we have \[r = - 6\left( \frac{ y }{ r } \right)\] can you finish it off?

OpenStudy (anonymous):

I'm not sure I understand where to go from there

OpenStudy (astrophysics):

\[r^2 = - 6y\]

OpenStudy (astrophysics):

What's next?

OpenStudy (anonymous):

square - 6y?

OpenStudy (astrophysics):

Why? Look at what we know, and we want to "get rid" of the polar coordinates.

OpenStudy (anonymous):

x^2?

OpenStudy (astrophysics):

I don't know what that means

OpenStudy (anonymous):

add it in?

OpenStudy (astrophysics):

Hint: \[r^2 = x^2+y^2\]

OpenStudy (anonymous):

-6y=x^2+y^2?

OpenStudy (astrophysics):

Yes, that looks good :)

OpenStudy (astrophysics):

You can rearrange it and what not if you wish

OpenStudy (anonymous):

That is it in terms of x and y?

OpenStudy (anonymous):

I'm asked to complete the square to produce another equation in my worksheet. Should I do it from this form?

OpenStudy (astrophysics):

You may complete the square

OpenStudy (astrophysics):

Ah, yes we have to complete the square, you know how to do that right?

OpenStudy (astrophysics):

When you have it as such \[-6y=x^2+y^2 \] it's always best to complete the square as it will be in terms of x and y.

OpenStudy (anonymous):

I know how to do it in regular form, but I'm sort of confused about this one

OpenStudy (astrophysics):

\[-6y = x^2+y^2 \implies x^2+y^2 + 6y = 0\]

OpenStudy (anonymous):

Oh, I thought it was when you added (b/2)^2 to both sides or something like that

OpenStudy (astrophysics):

I didn't complete the square...I put in a form so you can complete the square.

OpenStudy (anonymous):

Oh, okay gotcha

OpenStudy (astrophysics):

So we don't need to worry about the x^2, now complete the square for y^2+6y

OpenStudy (astrophysics):

When you have \[x^2+ax \implies x^2+ 2 \frac{ ax }{ 2 } + \left( \frac{ a }{ 2 } \right)^2-\left( \frac{ a }{ 2 } \right)^2\] to complete the square.

OpenStudy (anonymous):

Wouldn't that make it remain the same?

OpenStudy (anonymous):

Sorry, I suck at pre calc :(

OpenStudy (astrophysics):

You should try it, otherwise me doing everything is not going to help you :P

OpenStudy (anonymous):

very true

OpenStudy (anonymous):

But when plugging in y^2+6y in, it cancels back to y^2+6y, no?

OpenStudy (astrophysics):

Mhm, no you don't do that, we don't plug in anything I just put if you have the form y^2+6y you do the following -> .....etc

OpenStudy (astrophysics):

\[x^2+ax \implies x^2+ 2 \frac{ ax }{ 2 } + \left( \frac{ a }{ 2 } \right)^2-\left( \frac{ a }{ 2 } \right)^2\] \[x^2+ ax \implies y^2+6y\] in your question

OpenStudy (anonymous):

y ^2+6x+9-9?

OpenStudy (anonymous):

wait, no

OpenStudy (astrophysics):

 now we just factor